OpenStudy (anonymous):

the derivatives with respect to x of 2cos^2(x^2 + 2). a. 4sin(x^2 + 2) cos (x^2 + 2) b. -4sin (x^2 + 2)cos (x^2 + 2) c. 8xsin(x^2 + 2) cos (x^2 + 2) d. -8xsin(x^2 + 2) cos(x^2 + 2)

OpenStudy (anonymous):

d. -8xsin(x^2 + 2) cos(x^2 + 2)

sam (.sam.):

2cos^(2)(x^(2)+2) let (d)/(dv) cos(v)=-sin(v) (d)/(dx) x^(2)+2=2x (d)/(dx) 2cos^(2)(x^(2)+2)=-4xcos((x^(2)+2))sin((x^(2)+2)) -4xcos((x^(2)+2))sin((x^(2)+2))

sam (.sam.):

b

OpenStudy (anonymous):

sam that formula u use we use it now on calculus

OpenStudy (lalaly):

its d

OpenStudy (anonymous):

its D =D

OpenStudy (anonymous):

wew b or d? lets use b and d

sam (.sam.):

about the x

sam (.sam.):